OpenStudy (anonymous):

Which expressions are perfect square trinomials? Select all that apply. a.x^2+8x+12 b.x^2+5x+25/4 c.x^2+14x+29 d.x^2+7x+10.9

jhonyy9 (jhonyy9):

a.(x+2)(x+6)

jhonyy9 (jhonyy9):

b.(x+5/2)^2

TheSmartOne (thesmartone):

B is a perfect square trinomial. (x+5/2)^2

TheSmartOne (thesmartone):

the rest are not perfect square trinomials

OpenStudy (skullpatrol):

Any ideas @Tara_Kenny16 ?

OpenStudy (anonymous):

I was thinking it was c or d

OpenStudy (skullpatrol):

What do you know about perfect square trinomials?

OpenStudy (anonymous):

Nothing at the moment..

OpenStudy (skullpatrol):

A perfect square trinomial is an expression with three terms that is a square of a binomial.

OpenStudy (anonymous):

That makes sense.

OpenStudy (skullpatrol):

So there are two forms of binomials: a + b and a - b. $$(a+b)^2 = a^2 +2ab +b^2$$and$$(a-b)^2= a^2 -2ab+b^2$$